# 11footnotetext: Equal Contribution

### v1

@@ 1 Introduction
Figure 1: We plot the error rate for a hypothetical scenario modelling the high stakes exam described in Section 1. We model the error rate on each of the test portions as being proportional to $\propto\frac{1}{n_{i}^{\alpha}}$, where $n_{i}$ represents the studying budget spent on that portion of the exam, so $i=1$ corresponds to European History and $i=2$ to the Chinese History and set $n_{1}+n_{2}=N$ to be the total studying budget, with $N=100$ hours. The exponent $\alpha$ is $\alpha=1$ on the left plot and $\alpha=2$ on the right plot. In both cases, we consider $n_{1}=qN$ and $n_{2}=(1-q)N$, where $q$ is the proportion of time spent studying for the European History portion of the exam. This way, the error rate on the exam can be written as a function of $q$ as $L(q)=0.9\frac{1}{(100q)^{\alpha}}+0.1\frac{1}{(100q)^{\alpha}}$. We can see on both plots that shifting away from the testing proportion (red line, i.e. $q=90\%$) can lead to a better error rate with the optimal test proportion (green line, i.e., $q^{*}$ whose values are displayed accordingly). See also Corollary 3.3.
@@
In this paper, we focus on how positive distribution shift can arise purely due to mixture effects, and not because of “transfer” between components or components being more or less informative or useful. Positive distribution shift can certainly arrise also for other reason: E.g., it might be better to train on cleaner or less noisy data, or more generally leveraging transfer patterns between tasks or components Albalak et al. (2023); Liu et al. (2025); Jiang et al. (2025); Shukor et al. (2025). An even stronger benefit might be computational, where changes in the training distribution provide structure that is easier to exploit computationally, as is hinted by e.g. Abbe et al. (2023); Wang et al. (2025). All of these effects can be compounded with the mixture effect we study here. Indeed, several recent empirical papers looked at optimizing training mixture proportions in order to obtain good performance on a mixutre distribution such as Pile, showing empirically that the optimal training proportions differ from the test proportions Xie et al. (2023a); Ye et al. (2025); Albalak et al. (2023); Jiang et al. (2025); Shukor et al. (2025). As also indicated by the empirical scaling laws uncovered, this is due to a large part due to complex non-symetric transfer paterns between the different components, which could give rise to arbitrary Positive Distribution Shift paterns. But in this paper, we focus on understanding and mathematically charactarizing the direct effect of changing the mixture proportions, both as an important effect in their own right, and to better disentangle them from other effects when understanding Positive Distribution Shift forces in more complex problems involving also transfer and computational aspects. The papers mentioned here, as well as others González and Abu-Mostafa (2015); Hoffmann et al. (2022); Sorscher et al. (2022); Xie et al. (2023b); Gu et al. (2025), also emphasize the prevelance of Positive Distribution Shift in practice and how such “data set selection” is an important part of contemporary machine learning—it would thus benefit us to better understand and charactarize how and why it can happen and obtain a framework and language for discussing it.
@@ 2 Setup
Our main motivation, and the main way to interpret our work, is addressing how could we have positive distribution shift in the mixture setting, i.e., how could $D_{\mathrm{train}}\neq D_{\mathrm{test}}$ be better than $D_{\mathrm{train}}=D_{\mathrm{test}}$, and what qualitative changes in $D_{\mathrm{train}}$ make it better. This provides guidance in understanding positive distribution shift and seeking good training distributions. Nevertheless, there are also several realistic examples where it is conceivable the test mixing proportions are known and the training mixing proportions can be controlled or specified. For example, if we are pretraining a large language model for a set of tasks, we might well know their frequency at test time. Even in settings where the exact mixing proportions are unknown, it suffices to estimate the mixing proportions roughly by a quick analysis of test samples. If we suspect a mixture structure in the data, we can try to build crude classifiers for the components. This can be easy for a variety of tasks. For example, in a language-related task if each mixture component is a different language, or in a memorization task, if each mixture component is a topic or area (e.g., sports, science, etc.), we can build crude topic classifiers based on a small amount of data for each topic, or perhaps unsupervised clustering of a sample, then classify a sample. We can estimate the unknown mixing proportions using this classifier. Note that the classifier does not have to be very accurate since we do not care about individual errors, just about the aggregate proportions, and a bit of an error on the proportions is fine. Similar approach is taken in Ye et al. (2025), where the authors propose mixture dependent scaling laws for finding good training mixture proportion and assume that the validation data comes from an unknown mixture, which they estimate as part of their procedure in finding the good mixture proportions.
@@ 3 Power Law
Many machine learning tasks can be captured with power law error functions. Some classic examples include linear regression or learning VC classes, both of which have error rate $\propto\frac{1}{N}$, where $N$ is the number of data samples (Shalev-Shwartz and Ben-David, 2014). More recently, there have been many papers studying the loss curves of large language models for different tasks as a function of the compute budget or the number of training tokens through various scaling laws, which model error as having power law dependence on the number of data samples, such as the Chinchilla Scaling Law (Hoffmann et al., 2022), and many others (Kaplan et al., 2020; Cherti et al., 2023; Ye et al., 2025).
@@
In Theorem 3.2, we characterize the test error improvement from the positive distribution shift from optimal data mixing ratios in 3.1 when the size of the training data $N$ is large.
@@
In 3.1, if for the exponents it holds that $\alpha_{1}=\alpha_{2}=\dots=\alpha_{S}<\alpha_{S+1}\leq\alpha_{S+2}\leq\dots\leq\alpha_{K}$ for some $S$, then there exist $\varepsilon_{1},\varepsilon_{2}>0$ such that for any test data mixing ratio $\bm{p}$} and any $N>N_{0}(\{A_{i},B_{i},\alpha_{i},p_{i}\}_{i=1}^{K})$ we have that the following holds
@@
Theorem 3.2 shows that in the Power Law 3.1, positive distribution shift from optimal data mixing ratios improves the prefactor of the test error dependence on the number of data samples $N$ but does not change the decay rate in terms of $N$. For the proof of Theorem 3.2 and a more precise statement, including the closed form of $N_{0}(\{A_{i},B_{i},\alpha_{i},p_{i}\}_{i=1}^{K})$, see Section A.1.
@@
Consider 3.1 with $S=K$, i.e. $\alpha_{1}=\dots=\alpha_{K}=\alpha$, $A_{1}=\dots=A_{K}=A$, and $B_{1}=\dots=B_{k}=B$ with ${\bm{p}}=(p,\frac{1-p}{K-1},\dots,\frac{1-p}{K-1})$. We have that for any $\epsilon>0$
@@
Figure 2: We consider the setup of Corollary 3.3 with $A=1,~\alpha=0.28$, $K=100$, and some fixed NN. On the left plot, we show the “non-shifted” expected population loss $L^{\mathrm{same}}({\bm{p}})$ and the optimally mixed expected population loss $L^{*}({\bm{p}})$ as a function of majority population mass pp. On the right plot, we show the ratio of sample complexities for any fixed $\epsilon>0$, $N^{\textrm{ratio}}_{\epsilon}({\bm{p}})$ as a function of the mass of the majority population, pp. We can see significant improvement in the sample complexity from the positive distribution shift from using optimal mixing ratio, even up to $\approx 25\%$.
Corollary 3.3 demonstrates that if we have one majority population and a number of minority populations, the positive distribution shift from optimal data mixing ratio significantly improves sample complexity. For fixed $p$, if $K$ is large enough, $N^{\textrm{ratio}}({\bm{p}})$ will be close to $N^{\textrm{ratio}}({\bm{p}})\approx 1-p<1$, i.e. we get sample complexity improvement of up to $p$. For example, for $p=0.7$, $\alpha=0.28$, and $K=100$, for any $\epsilon>0$, $N_{\epsilon}^{\textrm{ratio}}({\bm{p}})\approx 0.75$, i.e. we achieve the same error with $\approx 25\%$ less data samples. We illustrate this in Figure 2. For the proof of Corollary 3.3, see Section A.1.
Furthermore, the test taking example considered in the introduction Section 1 follows from Corollary 3.3, by taking $K=2$, $\alpha=1$, and ${\bm{p}}=(0.9,0.1)$ (with any $A$ and taking $B$ much smaller than $1$). In particular, this shows that the optimal studying budget allocation is ${\bm{q}}^{*}=(0.75,0.25)$ and the improvement is $N^{\textrm{ratio}}({\bm{p}})=0.8$. This means that if you study for the exam with the right mixing ratio $\bm{q}$$^{*}$, you would need to study $20\%$ less time to achieve the same score as compared to using the test mixing ratio $\bm{p}$}. Further, taking $\alpha=\frac{1}{2}$ we get the second example on Figure 2. This shows that we indeed get ${\bm{q}}^{*}=(0.812\dots,0.188\dots)$ and $N^{\textrm{ratio}}({\bm{p}})=0.944$.
@@ 4 Memorization Tasks
The following theorem characterizes the test error improvement from the positive distribution shift from optimal data mixing ratios in the Memorization Task 4.1.
@@
In 4.1, for all ${\bm{p}}\in\Delta^{K-1}$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{K}$, the expected loss when training on $n$ samples is given by
@@
For example, when $\alpha=1.5$, we have $L^{\mathrm{same}}({\bm{p}})=\Theta(N^{-1/3})$ and $L^{*}({\bm{p}})=\Theta(N^{-1/2})$. For the proof of Corollary 4.3, see Section A.2.
@@ 6 Transfer Learning
In this section, we also consider a setting where the tasks are not independent, and there is transfer. Generally, with transfer and in multitask learning, the error on the $k$-th task is affected by the number of samples on other tasks, that is, the error on task $k$ decreases if we hold $n_{k}$ constant but increase the number of samples on related tasks. In the most typical transfer learning setups studied in the literature, such as multi-task learning of linear classifiers over linear representation with feature learning (Baxter, 2011; Maurer, 2009; Pontil and Maurer, 2013; Aliakbarpour et al., 2024) and multi-task learning with shared sparsity (Wang et al., 2016, 2017), the transfer effect is captured by the following model, with a slight extension of our framework.
@@
For example, in multi-task learning of shared sparsity (Wang et al., 2017), the error bound takes this form with $\alpha_{1}=\dots=\alpha_{K}=1$. Interestingly, it turns out that the behavior of the test error in these typical transfer learning settings is the same with respect to the mixture proportions as for the independent component setting, and so our analysis is actually applicable to this non-independent setting as well. The Standard Transfer Learning 6.1 is equivalent to the setup of Power Law Tasks 3.1 in the sense that we can understand the optimal data mixing ratio $\bm{q}$$^{*}$ and the error improvement of the Standard Transfer Learning model from a specific instance of the Power Law 3.1. Namely, the transfer term in each of the subpopulation loss functions can be decomposed into a transfer error term and a specific task error term $e_{k}({\bm{n}})=e_{k}^{\textrm{transfer}}({\bm{n}})+e_{k}^{\textrm{spec}}({\bm{n}})$, where $e_{k}^{\textrm{transfer}}({\bm{n}})=\frac{A_{0,k}}{(n_{1}+\dots+n_{k})^{\alpha_{k}}+B_{0,k}}$ is independent of the distribution of samples across different tasks, and $e_{k}^{\textrm{spec}}({\bm{n}})=e_{k}^{\textrm{spec}}(n_{k})=\frac{A_{1,k}}{n_{k}^{\alpha_{k}}+B_{1,k}}$ only depends on $n_{k}$. Therefore, the transfer error term $e_{k}^{\textrm{transfer}}({\bm{n}})$ in each of the subpoluation error functions will only offset the final expected loss $L({\bm{p}},{\bm{q}})$ by $\sum_{i=1}^{K}p_{i}\frac{A_{0,k}}{N^{\alpha_{k}}+B_{0,k}}$, which only depends on the total number of samples $N$. On the other hand, the specific task error terms $e_{k}^{\textrm{spec}}(n_{k})$ can be thought of as independent (i.e. without transfer) tasks and will behave the same as in 3.1. So, for the Standard Transfer Learning 6.1, the optimal data mixing ratio $\bm{q}$$^{*}$ and the expected test losses $L^{*}({\bm{p}})$ and $L^{\mathrm{same}}({\bm{p}})$ are given by Equation 1, Equation 2 and Theorem 3.2 respectively in Theorem 3.2 with $A_{k}$ being replaced by $A_{1,k}$.
@@ 7 It’s Almost Always Better to Mismatch
More precisely, we will show that if there is no positive distribution shift, either the test data mixing ratio is on a measure zero set of the simplex or the subpopulation error functions $e_{k}(n_{k})$ have to be all constant functions, which is meaningless. We show this in Corollary 7.4.
@@
For any set of subpopulations ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{K}$ and any learning algorithm $\mathcal{A}$}, either 7.1 holds, or there exists a zero-measure set $U$ on $\Delta_{K-1}$ such that for all ${\bm{p}}\in\Delta_{K-1}\setminus U$, $L_{N}^{*}({\bm{p}})<L^{\mathrm{same}}({\bm{p}})$.
Theorem 7.2 shows that either $\bm{p}$} is on a measure zero set $U$ on $\Delta_{K-1}$ or the Conservation 7.1 must hold. Next, we show that if the tasks are independent, then the Conservation 7.1 holds only if all of the subpopulation error functions are constants.
@@
Theorem 7.2 and Lemma 7.3 together show that positive distirbution shift always exists, unless all the subpopulation error functions are constant.
@@
For the proofs of Theorem 7.2, Lemma 7.3, and Corollary 7.4, see Appendix B.

### v2

# Shift is Good: Mismatched Data Mixing Improves Test Performance

@@ 1 Introduction
Figure 1: We plot the error rate for a hypothetical scenario modelling the high stakes exam described in Section 1. We model the error rate on each of the test portions as being proportional to $\propto\frac{1}{n_{i}^{\alpha}}$, where $n_{i}$ represents the studying budget spent on that portion of the exam, so $i=1$ corresponds to European History and $i=2$ to the Chinese History and set $n_{1}+n_{2}=N$ to be the total studying budget, with $N=100$ hours. The exponent $\alpha$ is $\alpha=1$ on the left plot and $\alpha=2$ on the right plot. In both cases, we consider $n_{1}=qN$ and $n_{2}=(1-q)N$, where $q$ is the proportion of time spent studying for the European History portion of the exam. This way, the error rate on the exam can be written as a function of $q$ as $L(q)=0.9\frac{1}{(100q)^{\alpha}}+0.1\frac{1}{(100q)^{\alpha}}$. We can see on both plots that shifting away from the testing proportion (red line, i.e. $q=90\%$) can lead to a better error rate with the optimal test proportion (green line, i.e., $q^{*}$ whose values are displayed accordingly). See also Section 3.
@@
In this paper, we focus on how positive distribution shift can arise purely due to mixture effects, and not because of “transfer” between components or components being more or less informative or useful. Positive distribution shift can certainly arrise also for other reason: E.g., it might be better to train on cleaner or less noisy data, or more generally leveraging transfer patterns between tasks or components albalak2023efficientonlinedatamixing; liu2025regmix; jiang2025adaptive; shukor2025scalinglawmixing. An even stronger benefit might be computational, where changes in the training distribution provide structure that is easier to exploit computationally, as is hinted by e.g. abbe2023provable; wang2025learningcompositionalfunctionstransformers. All of these effects can be compounded with the mixture effect we study here. Indeed, several recent empirical papers looked at optimizing training mixture proportions in order to obtain good performance on a mixutre distribution such as Pile, showing empirically that the optimal training proportions differ from the test proportions xie2023doremioptimizingdatamixtures; ye2025mixing; albalak2023efficientonlinedatamixing; jiang2025adaptive; shukor2025scalinglawmixing. As also indicated by the empirical scaling laws uncovered, this is due to a large part due to complex non-symetric transfer paterns between the different components, which could give rise to arbitrary Positive Distribution Shift paterns. But in this paper, we focus on understanding and mathematically charactarizing the direct effect of changing the mixture proportions, both as an important effect in their own right, and to better disentangle them from other effects when understanding Positive Distribution Shift forces in more complex problems involving also transfer and computational aspects. The papers mentioned here, as well as others gonzalez2015mismatched; hoffmann2022computeoptimal; sorscher2022beyond; xie2023data; gu2025data, also emphasize the prevelance of Positive Distribution Shift in practice and how such “data set selection” is an important part of contemporary machine learning—it would thus benefit us to better understand and charactarize how and why it can happen and obtain a framework and language for discussing it.
@@ 2 Setup
Our main motivation, and the main way to interpret our work, is addressing how could we have positive distribution shift in the mixture setting, i.e., how could $D_{\mathrm{train}}\neq D_{\mathrm{test}}$ be better than $D_{\mathrm{train}}=D_{\mathrm{test}}$, and what qualitative changes in $D_{\mathrm{train}}$ make it better. This provides guidance in understanding positive distribution shift and seeking good training distributions. Nevertheless, there are also several realistic examples where it is conceivable the test mixing proportions are known and the training mixing proportions can be controlled or specified. For example, if we are pretraining a large language model for a set of tasks, we might well know their frequency at test time. Even in settings where the exact mixing proportions are unknown, it suffices to estimate the mixing proportions roughly by a quick analysis of test samples. If we suspect a mixture structure in the data, we can try to build crude classifiers for the components. This can be easy for a variety of tasks. For example, in a language-related task if each mixture component is a different language, or in a memorization task, if each mixture component is a topic or area (e.g., sports, science, etc.), we can build crude topic classifiers based on a small amount of data for each topic, or perhaps unsupervised clustering of a sample, then classify a sample. We can estimate the unknown mixing proportions using this classifier. Note that the classifier does not have to be very accurate since we do not care about individual errors, just about the aggregate proportions, and a bit of an error on the proportions is fine. Similar approach is taken in ye2025mixing, where the authors propose mixture dependent scaling laws for finding good training mixture proportion and assume that the validation data comes from an unknown mixture, which they estimate as part of their procedure in finding the good mixture proportions.
@@ 3 Power Law
Many machine learning tasks can be captured with power law error functions. Some classic examples include linear regression or learning VC classes, both of which have error rate $\propto\frac{1}{N}$, where $N$ is the number of data samples (shalev2014understanding). More recently, there have been many papers studying the loss curves of large language models for different tasks as a function of the compute budget or the number of training tokens through various scaling laws, which model error as having power law dependence on the number of data samples, such as the Chinchilla Scaling Law (hoffmann2022computeoptimal), and many others (kaplan2020scalinglawsneurallanguage; cherti2023scaling; ye2025mixing).
@@
In Theorem 3.2, we characterize the test error improvement from the positive distribution shift from optimal data mixing ratios in Section 3 when the size of the training data $N$ is large.
@@
In Section 3, if for the exponents it holds that $\alpha_{1}=\alpha_{2}=\dots=\alpha_{S}<\alpha_{S+1}\leq\alpha_{S+2}\leq\dots\leq\alpha_{K}$ for some $S$, then there exist $\varepsilon_{1},\varepsilon_{2}>0$ such that for any test data mixing ratio $\bm{p}$} and any $N>N_{0}(\{A_{i},B_{i},\alpha_{i},p_{i}\}_{i=1}^{K})$ we have that the following holds
@@
Theorem 3.2 shows that in the Power Law Section 3, positive distribution shift from optimal data mixing ratios improves the prefactor of the test error dependence on the number of data samples $N$ but does not change the decay rate in terms of $N$. For the proof of Theorem 3.2 and a more precise statement, including the closed form of $N_{0}(\{A_{i},B_{i},\alpha_{i},p_{i}\}_{i=1}^{K})$, see Section A.1.
@@
Consider Section 3 with $S=K$, i.e. $\alpha_{1}=\dots=\alpha_{K}=\alpha$, $A_{1}=\dots=A_{K}=A$, and $B_{1}=\dots=B_{k}=B$ with ${\bm{p}}=(p,\frac{1-p}{K-1},\dots,\frac{1-p}{K-1})$. We have that for any $\epsilon>0$
@@
Figure 2: We consider the setup of Section 3 with $A=1,~\alpha=0.28$, $K=100$, and some fixed NN. On the left plot, we show the “non-shifted” expected population loss $L^{\mathrm{same}}({\bm{p}})$ and the optimally mixed expected population loss $L^{*}({\bm{p}})$ as a function of majority population mass pp. On the right plot, we show the ratio of sample complexities for any fixed $\epsilon>0$, $N^{\textrm{ratio}}_{\epsilon}({\bm{p}})$ as a function of the mass of the majority population, pp. We can see significant improvement in the sample complexity from the positive distribution shift from using optimal mixing ratio, even up to $\approx 25\%$.
Section 3 demonstrates that if we have one majority population and a number of minority populations, the positive distribution shift from optimal data mixing ratio significantly improves sample complexity. For fixed $p$, if $K$ is large enough, $N^{\textrm{ratio}}({\bm{p}})$ will be close to $N^{\textrm{ratio}}({\bm{p}})\approx 1-p<1$, i.e. we get sample complexity improvement of up to $p$. For example, for $p=0.7$, $\alpha=0.28$, and $K=100$, for any $\epsilon>0$, $N_{\epsilon}^{\textrm{ratio}}({\bm{p}})\approx 0.75$, i.e. we achieve the same error with $\approx 25\%$ less data samples. We illustrate this in Figure 2. For the proof of Section 3, see Section A.1.
Furthermore, the test taking example considered in the introduction Section 1 follows from Section 3, by taking $K=2$, $\alpha=1$, and ${\bm{p}}=(0.9,0.1)$ (with any $A$ and taking $B$ much smaller than $1$). In particular, this shows that the optimal studying budget allocation is ${\bm{q}}^{*}=(0.75,0.25)$ and the improvement is $N^{\textrm{ratio}}({\bm{p}})=0.8$. This means that if you study for the exam with the right mixing ratio $\bm{q}$$^{*}$, you would need to study $20\%$ less time to achieve the same score as compared to using the test mixing ratio $\bm{p}$}. Further, taking $\alpha=\frac{1}{2}$ we get the second example on Figure 2. This shows that we indeed get ${\bm{q}}^{*}=(0.812\dots,0.188\dots)$ and $N^{\textrm{ratio}}({\bm{p}})=0.944$.
@@ 4 Memorization Tasks
The following theorem characterizes the test error improvement from the positive distribution shift from optimal data mixing ratios in the Memorization Task Section 4.
@@
In Section 4, for all ${\bm{p}}\in\Delta^{K-1}$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{K}$, the expected loss when training on $n$ samples is given by
@@
For example, when $\alpha=1.5$, we have $L^{\mathrm{same}}({\bm{p}})=\Theta(N^{-1/3})$ and $L^{*}({\bm{p}})=\Theta(N^{-1/2})$. For the proof of Section 4, see Section A.2.
@@ 6 Transfer Learning
In this section, we also consider a setting where the tasks are not independent, and there is transfer. Generally, with transfer and in multitask learning, the error on the $k$-th task is affected by the number of samples on other tasks, that is, the error on task $k$ decreases if we hold $n_{k}$ constant but increase the number of samples on related tasks. In the most typical transfer learning setups studied in the literature, such as multi-task learning of linear classifiers over linear representation with feature learning (baxter2011inductive; Maurer2009TransferBF; pontil2013excess; aliakbarpour2024metalearning) and multi-task learning with shared sparsity (wang2016distributed; wang2017distributed), the transfer effect is captured by the following model, with a slight extension of our framework.
@@
For example, in multi-task learning of shared sparsity (wang2017distributed), the error bound takes this form with $\alpha_{1}=\dots=\alpha_{K}=1$. Interestingly, it turns out that the behavior of the test error in these typical transfer learning settings is the same with respect to the mixture proportions as for the independent component setting, and so our analysis is actually applicable to this non-independent setting as well. The Standard Transfer Learning Section 6 is equivalent to the setup of Power Law Tasks Section 3 in the sense that we can understand the optimal data mixing ratio $\bm{q}$$^{*}$ and the error improvement of the Standard Transfer Learning model from a specific instance of the Power Law Section 3. Namely, the transfer term in each of the subpopulation loss functions can be decomposed into a transfer error term and a specific task error term $e_{k}({\bm{n}})=e_{k}^{\textrm{transfer}}({\bm{n}})+e_{k}^{\textrm{spec}}({\bm{n}})$, where $e_{k}^{\textrm{transfer}}({\bm{n}})=\frac{A_{0,k}}{(n_{1}+\dots+n_{k})^{\alpha_{k}}+B_{0,k}}$ is independent of the distribution of samples across different tasks, and $e_{k}^{\textrm{spec}}({\bm{n}})=e_{k}^{\textrm{spec}}(n_{k})=\frac{A_{1,k}}{n_{k}^{\alpha_{k}}+B_{1,k}}$ only depends on $n_{k}$. Therefore, the transfer error term $e_{k}^{\textrm{transfer}}({\bm{n}})$ in each of the subpoluation error functions will only offset the final expected loss $L({\bm{p}},{\bm{q}})$ by $\sum_{i=1}^{K}p_{i}\frac{A_{0,k}}{N^{\alpha_{k}}+B_{0,k}}$, which only depends on the total number of samples $N$. On the other hand, the specific task error terms $e_{k}^{\textrm{spec}}(n_{k})$ can be thought of as independent (i.e. without transfer) tasks and will behave the same as in Section 3. So, for the Standard Transfer Learning Section 6, the optimal data mixing ratio $\bm{q}$$^{*}$ and the expected test losses $L^{*}({\bm{p}})$ and $L^{\mathrm{same}}({\bm{p}})$ are given by Equation 1, Equation 2 and Theorem 3.2 respectively in Theorem 3.2 with $A_{k}$ being replaced by $A_{1,k}$.
@@ 7 It’s Almost Always Better to Mismatch
More precisely, we will show that if there is no positive distribution shift, either the test data mixing ratio is on a measure zero set of the simplex or the subpopulation error functions $e_{k}(n_{k})$ have to be all constant functions, which is meaningless. We show this in Section 7.
@@
For any set of subpopulations ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{K}$ and any learning algorithm $\mathcal{A}$}, either Section 7 holds, or there exists a zero-measure set $U$ on $\Delta_{K-1}$ such that for all ${\bm{p}}\in\Delta_{K-1}\setminus U$, $L_{N}^{*}({\bm{p}})<L^{\mathrm{same}}({\bm{p}})$.
Theorem 7.2 shows that either $\bm{p}$} is on a measure zero set $U$ on $\Delta_{K-1}$ or the Conservation Section 7 must hold. Next, we show that if the tasks are independent, then the Conservation Section 7 holds only if all of the subpopulation error functions are constants.
@@
Theorem 7.2 and Section 7 together show that positive distirbution shift always exists, unless all the subpopulation error functions are constant.
@@
For the proofs of Theorem 7.2, Section 7, and Section 7, see Appendix B.